40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Nov 5 19:38:52 2012, jamie, jenne, ayaka, den, Update, Alignment, more alignment woes 
    Reply  Tue Nov 6 16:38:37 2012, jenne, jamie, ayaka, manasa, Update, Alignment, Alignment back under control again 
       Reply  Tue Nov 6 17:07:04 2012, jamie, Update, Alignment, AS and REFL tmp6oTENk.pngtmplamEtZ.png
       Reply  Tue Nov 6 17:22:51 2012, Manasa, Jamie, Update, Alignment, Alignment- POY and oplevs 
          Reply  Wed Nov 7 00:10:38 2012, Jenne , Update, Alignment, Alignment- POY and oplevs. photos. 
             Reply  Wed Nov 7 07:11:10 2012, rana, Update, Alignment, Alignment- POY and oplevs. photos. 
             Reply  Wed Nov 7 09:09:02 2012, Steve, Update, Alignment, Alignment- 9amNov7w.png
                Reply  Wed Nov 7 15:17:15 2012, Steve, Update, Alignment, steering option with pico motor? IMG_1795.JPG
Message ID: 7673     Entry time: Tue Nov 6 16:38:37 2012     In reply to: 7671     Reply to this: 7674   7675
Author: jenne, jamie, ayaka, manasa 
Type: Update 
Category: Alignment 
Subject: Alignment back under control again 

We had a big alignment party early this morning, and things are back to looking good.  We have been very careful not to bump or touch tables any more than necessary.  Also, we have removed the apertures from the BS and PRM, so there are no more apertures currently left in the chambers (this is good, since we won't forget).

We started over again from the PZTs, using the PRM aperture and the freestanding aperture in front of PR2, to get the height of the beam correct.  We then moved PZTs to get the beam centered on BS, ITMY, ETMY.  We had to do a little poking of PR2 (and PR3?) to get pitch correct everywhere.

We then went to ETMX to check beam pointing, and used BS to steer the beam to the center of ETMX.  We checked that the beam was centered on ITMX.

We went through and ensured that ITMX, ITMY, PRM, SRM are all retroreflecting.  We see nice MICH fringes, and we see some fringes (although still not so nice...) when we bring PRM and SRM into alignment.

We checked the AS path (with only MICH aligned), and made sure we are centered on all of the mirrors.  This included steering a little bit on the mirrors on the OMC table, in yaw.  Initially, AS was coming out of the vacuum, but hitting the side of the black beam tube.  Now it gets nicely to the table.

For both AS and REFL, we made sure there is no clipping in the OMC chamber.

I recentered the beams for AS and REFL on their respective cameras.

IPPOS was centered on the QPD.  This involved moving the first out-of-vac steering mirror sideways a small amount, since the beam was hitting the edge of the mirror.  IPANG was aligned in-vac, and has been centered on the QPD.

Right now, Manasa, Jamie and Ayaka are doing some finishing touches work, checking that POY isn't clipping on OM2, the second steering mirror after the SRM, and they'll confirm that POX comes out of the chamber nicely, and that POP is also still coming out (by putting the green laser pointer back on that table, and making sure the green beam is co-aligned with the beam from PR2-PR3.  Also on the list is checking the vertex oplevs.  Steve and Manasa did some stuff with the ETM oplevs yesterday, but haven't had a chance to write about it yet.

ELOG V3.1.3-